[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Wed Jul 19, 2017

Monica Bozeman - NOAA Federal monica.bozeman at noaa.gov
Wed Jul 19 20:15:23 UTC 2017


Dan,

THe Himawari data was turned off to prevent the satellite data partition
filling up until we upgrade to the new netapp next week.

Monica


Monica Bozeman
Meteorologist/Programmer
National Hurricane Center
301-683-1553
monica.bozeman at noaa.gov
<http://www.wdtb.noaa.gov/badges/monica.bozeman>

On Wed, Jul 19, 2017 at 8:13 PM, HSU Shift Log <nhc.tsbadmin at noaa.gov>
wrote:

> *HSU Shift Log for Wed Jul 19 2017 16:13:14 GMT-0400 (EDT)*
>
> *Shift:* H
>
> *Forecaster:* D Brown
>
> *Technical Issues Summary:*
>
> *ATCF issues:*
>
> *N-AWIPS issues:*
>
> *AWIPS-2/NCP issues:*
>
> *Other Techincal issues:* Himawari satellite data is not available in
> NAWIPS.
>
> *Miscellaneous issues:* Strong smell of electrical burning developed in
> operations around 1:15 pm. Neal was notified and began to detect the
> source. The smell became stronger and Jann, Patty, Dennis, and the guard
> began enacting emergency plans. The source of the burning could not be
> found and it was decided to sound the fire alarm, evacuate the building,
> and call the fire department. The TWO were ready and sent for 18z. Hugh
> Cobb notified the SDM and Mike contacted CPHC about potential backup. The
> fire department gave the all-clear around 1:45 pm, and everyone was allowed
> back into the building. It did not affect our ability to issue the 21z
> packages. CPHC was notified.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170719/1bf484e5/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list